Supplementary Information for PEC for Doron ZEILBERGER

Last Update: Jan. 27, 2010

Honors

1. I have been invited to give the American Mathematical Society’s 2010 Erdös Memorial Lecture. This annual lecture is considered a great honor (previous speakers include a Fields medalist (Gowers), a Wolf prize winner (Furstenberg), a Nevallina prize-winner (Wigderson) a MacArthur fellow (Diaconis) and many other luminaries (Bollobas, Conway, Graham, Lagarias, Sturmfels).

2. I have been invited to give the 2010 Chaim Leib Pekeris lecture at the Weizmann Institute of Science in Israel. This annual lecture is considered a great honor (previous speakers include a Nobel-prize winner (Aumann), two Fields-medalists (Smale, Zelmanov), three Wolf-prize winners (Langlands, Lovasz, Smale), two Turing-award [The “Nobel” of computer science] winners (Pnueli, Karp), three MacArthur fellows (E. Keller, Koller, Oster), two recipients of the National Medal of Science (Karp, Nirenberg), one Kyoto-prize winner (Karp), one knight (Sir James Lighthill) and other luminaries (Sarnak, Alon, Papadimitriou, …).

Grant

I was granted a five-year NSF grant for 2009-2014 totalling, so far, $350,000. It will likely have additional funds via future supplement to support graduate students, like in previous grants.

Graduate Education

1. My Experimental Mathematics graduate class is a regular feature, given every Spring semester, and is a huge success. The content changes each time, but students gain a solid foundation in computer algebra. I currently have 13 students, which, as you know, is very high for a nonrequirement graduate class.

2. Two students got their PhD, under my supervision, last academic year. Eric Rowland, currently a visting assistant professor at Tulane, and Paul Raff, who is now a postdoctoral fellow in the school of communication here at Rutgers.

3. I currently have three PhD students under my supervision: Andrew Baxter, Emilie Hogan, and Brian Nakamura. Their progress is very satisfactory.

Undergraduate Education

I taught two large lectures (each with more than 110 students) in multi-variable calculus (Math 251). There is a lot of teaching notes in my website: http://www.math.rutgers.edu/ Tilde zeilberg/DrZhandounts2nd.html, available for the whole world.

In the teaching evaluations, on the question “rate the teaching effectiveness” I got 4.87 and 4.76 (out of 5) in the these classes.

Seminar

I organize (with grad student Andrew Baxter) the weekly Experimental Mathematical seminar that is in its seventh consecutive year. We often have excellent out-of-town speakers.